# Large Language Models Play StarCraft II: Benchmarks and A Chain of Summarization Approach

### v1

@@ 5.1 Evaluation of the Chain of Summarization Method
5.2 Impact of Different Prompts on Strategy Formulation
Table 2: Comparison of LLM Agents’ Win Rates Against Various Difficulty Levels of Built-in AI in TextStarCraft II with Chain of Summarization Using Different Prompts. Notably, Prompt2 significantly enhances LLM agent performance, enabling victories against the Harder difficulty level of built-in AI.
LLMs Difficulty Level
VeryEasy Easy Medium Hard Harder Very Hard
GPT3.5-turbo-16k with prompt1 2/2 TBD TBD 0/2 TBD TBD
GPT3.5-turbo-16k with prompt2 8/8 9/9 8/8 21/25 7/14 0/12
We used the GPT3.5-turbo-16k model as our LLM, applying the Chain of Summarization method with two different types of prompts. In this context, ’prompt’ specifically refers to the prompts used in the multi-frame summarization method. We evaluated the effectiveness of these prompts in the TextStarCraft II environment, where the LLMs, playing as Protoss, faced off against Zerg opponents of varying difficulties. The results, displayed in Table 2, demonstrate that Prompt2 significantly enhances the performance of the LLM agents.
Simple Thought Chain :
In this context, we employ prompt1(In appendix Prompt1) to initiate a fundamental reasoning sequence, enabling the LLM agent to perform an elementary understanding and analysis of StarCraft II dynamics. This approach facilitates basic operations like generating workers, establishing new bases, producing combat units (e.g., Protoss Zealots or Zerg Roaches), and conducting reconnaissance. Nevertheless, it falls short of undertaking research upgrades or accessing advanced technologies, signifying that prompt1 limits the LLM’s capacity for in-depth strategy formulation or comprehensive awareness of the ongoing game scenario.
Complex Thought Chain :
Prompt2 (In appendix Prompt2) assimilates insights regarding StarCraft II’s intricate mechanics, necessitating that the LLM navigates six critical phases: Situation Overview, Situation Analysis, Strategic Planning, Opponent Strategy Analysis, Strategic Recommendations, and the Decision-Making Process(In appendix Prompt2). This holistic cognitive pathway empowers the LLM agent to engage in nuanced game aspects, such as initiating research upgrades, exploring the tech tree, erecting defensive infrastructures, transitioning military compositions, and orchestrating diversified forces. Demonstrably, it exhibits proficiency in clinching victories over the "buil-in" AI,"Harder"(level 5) difficulty level opponents, underlining its strategic acumen in complex gaming scenarios.
We discovered that it’s necessary to teach LLMs how to analyze the StarCraft II game. Formerly simple analytical prompts do not suffice for the LLM to understand the specialized information within StarCraft II. We need to design prompts reflecting the intricate thought processes of StarCraft II players, based on the game’s characteristics, to enhance LLMs’ comprehension abilities regarding StarCraft II itself.
5.3 Performance Evaluation of Various Large Language Models
Here, we evaluated only GPT3.5-Turbo, GPT3.5-Turbo-16k, GPT4, and models finetuned with data collected via GPT, namely Finetune-ChatGlm2 6b and Finetune-Llama2 7b. When testing the capabilities of different Large Language Models (LLMs), we uniformly applied the chain of summarization method and consistent prompts to assess their performance in the TextStarCraft II environment(Figure 1). In this context, we set LLMs to function as Protoss players, contending against built-in Zerg opponents of Harder.
GPT3.5:
We observed performance differences between GPT3.5-Turbo and GPT3.5-Turbo-16k within the TextStarCraft II scenarios. Our TextStarCraft II setting is highly complex, and under the chain of summarization approach, it requires the LLM to undertake intricate analysis and understanding of multi-frame information. The GPT3.5-Turbo model, with its 4k window, could generate only the analysis segment due to window size constraints, failing to produce the final suggestion and decision sections, resulting in an inability to play TextStarCraft II effectively. In contrast, the 16k version, with its extended processing window, does not truncate prematurely and can provide complete outputs. In extensive testing under these conditions, we found that the LLM agent could defeat AI opponents up to the "Harder" difficulty.
GPT4:
We also conducted limited testing with GPT4. We utilized GPT4 for tests against the built-in "Harder" difficulty within TextStarCraft II. Our tests revealed that GPT4 is incapable of overcoming the built-in AI at the "Very Hard" level. Observations from replaying the matches indicated that GPT4’s performance did not significantly surpass that of GPT3.5 16k.
Finetune-ChatGlm2 6b
[5] and Finetune-Llama2 7b[21]: We employed the GPT3.5-Turbo-16k model as the LLM agent, using data from interactions under TextStarCraft II to finetune these two models. We discovered that, within the complex TextStarCraft II environment, even with game-specific data finetuning, these models could not comprehend the intricacies of StarCraft II gameplay, producing disorganized responses. They remained incapable of effectively playing TextStarCraft II.
5.4 Assessing LLMs’ Mastery of StarCraft II Concepts
(a) GPT4 Evaluation of Questions
(b) GPT4 Evaluation of LLMs
(c) Human Evaluation of Questions
(d) Human Evaluation of LLMs
Figure 3: The result of Double-Blind Assessment.In our double-blind assessment, 30 human experts were invited to conduct evaluations, and separately, GPT-4 was used to perform 30 evaluations. The results revealed that Bard’s responses were the least effective among the LLMs, particularly in questions 3, 4, and 5. It was observed that GPT-4 tended to give higher scores to GPT-3.5 and itself, indicating a significant variance in ratings across different LLMs. In contrast, human experts, while rating GPT-4 highly, generally scored the other models more evenly, demonstrating less disparity in their evaluations.
Preliminary Evaluation:
After the release of GPT-4, we utilized build order data from renowned StarCraft II forums such as Team Liquid [^11] https://liquipedia.net/starcraft2/Main_Page and Spawning Tool [^12] https://lotv.spawningtool.com/ to assess the understanding of StarCraft II by GPT-3.5 and GPT-4. We observed that these models demonstrated a decent level of understanding and analytical ability regarding game situations. However, they often overlooked aspects like the tech tree and supply constraints, leading to shortcomings in accuracy.
In-depth Evaluation:
To explain why LLM agents can play StarCraft II using the CoS method in the TextStarCraft II environment, we propose that these LLMs have acquired extensive knowledge about StarCraft II during their pre-training. This encompasses extensive textual knowledge about StarCraft II, covering aspects like build orders, game details, the history of the game’s development, and esports information.To validate our hypothesis that LLM agents have acquired a wealth of StarCraft II knowledge during their pre-training, we adopted a unique approach. We posed questions to the language models and had human experts and GPT4 models blindly rate the responses. This method assessed the LLMs’ understanding of StarCraft II knowledge. The models evaluated were GPT-3.5, GPT-4, Claude2, and Bard. Our questions covered five key areas:
  1. 1.
    Basic Knowledge of StarCraft II.
  2. 2.
    Racial Mechanics in StarCraft II.
  3. 3.
    Typical Strategies in StarCraft II for Different matchups.
  4. 4.
    Standard build orders for each race.
  5. 5.
    Classic Strategies and Counterplays in StarCraft II.
For detailed questions and LLM responses, please refer to the appendixB.
Evaluation Methodology: We employed a double-blind evaluation method, hiding and randomizing the order of responses during the review and scoring by human experts and GPT-4. This approach mitigates biases to a certain extent, focusing solely on the content and comprehension of the responses themselves.
Evaluation Results: Our evaluations by 30 human experts and GPT-4 (see Figures 3) reveal that:
  • ChatGPT and GPT4: Both human experts and GPT-4 evaluations indicate that GPT-4 and ChatGPT perform significantly better than other models in addressing StarCraft II questions, with GPT-4 outperforming ChatGPT.Their responses align more closely with the game dynamics of StarCraft II and match players’ common understanding. Notably, GPT-4 can generate professional build ordersB.4 that take into account the game’s mechanics and unique racial features, rivaling the expertise found on professional gaming websites.
  • Bard: Bard’s responses in the StarCraft II QA were not as effective as those of other LLMs, especially in areas covered by questions 3, 4, and 5. Both human experts and GPT-4 evaluations highlighted significant errors in Bard’s answers, expressing dissatisfaction.Notably, Bard’s build orders did not align with the game’s dynamics and logic. This misalignment is a critical flaw for decision-making in StarCraft II, significantly impacting the effectiveness of gameplay strategies.
  • Claude2: Although Claude2’s performance is not as strong as GPT-4 and it shows considerable variability in answer quality, it generally satisfies human experts in most questions, achieving results comparable to ChatGPT.
Based on the combined ratings from human experts and GPT-4, the models’ proficiency in StarCraft II knowledge is ranked as follows: GPT-4 > GPT-3.5 > Claude2 > Bard. Despite varying levels of ability, all these models have acquired a substantial amount of StarCraft II-related knowledge through pre-training. Interestingly, when GPT-4 was involved in the evaluation, it tended to give higher scores to GPT-3.5 and itself, but lower scores to Bard. GPT-4’s ratings showed significant variances across different LLMs. In contrast, human experts, while also rating GPT-4 highly, tended to score the other LLMs more evenly, without such large disparities.
5.5 Strategic Patterns and Behavior Analysis of ChatGPT in StarCraft II
Figure 4: Unit preferences of starcraft II players
We conducted a further analysis of ChatGPT’s performance and strategic choices in StarCraft II, as illustrated in Figure 4. This analysis draws on data from 30 replays involving the GPT3.5 16K model, which employed the chain of summarization method and played as Protoss against the built-in Zerg AI on "Harder(lv5)" difficulty. Our study chiefly investigates ChatGPT’s army composition and unit preferences, uncovering pronounced patterns and specific strategic tendencies in its decision-making process.
Untapped Units: Despite having access to the necessary tech tree, LLM agents consistently avoid using several units, including the Mother ship, dark templar, archon, and warp prism. This suggests a significant divergence in strategy from top human players, possibly underestimating or overlooking the strategic importance of these units.
Infrequently Used Units: LLM agents show a cautious tendency in training combat units, rarely using oracle, adepts, carriers, tempests, disruptors, sentries, and high templars. This reflects some deficiencies in strategic diversity and adaptability compared to top human players, as these units usually require more intricate micromanagement, which may not be LLM Agents’ forte.
Commonly Used Units: LLM agents prefer using units such as observers, phoenixes, zealots, stalkers, colossi, void rays, and immortals. This indicates a more defense-oriented gameplay style compared to professional human players, favoring units that require less micromanagement in combat.
The strategy of LLM agents resembles that of Protoss players during the ’Heart of the Swarm’ era of StarCraft II(Figure 2). It focuses on economic development, building defenses, and upgrading technology, showcasing robust tactics commonly employed by skilled players. However, there is a significant gap between its strategy and that of professional human players, lacking creativity and exhibiting a more monotonous gameplay style."
5.6 Policy interpretability
Even the AI like AlphaStar, capable of defeating human professional players in StarCraft II, often makes irrational or inexplicable decisions and actions. In contrast, while the LLM agent may not excel at micromanagement like AlphaStar, its capacity for thinking and reasoning enables it to analyze and predict, thereby making more rational decisions.We summrize our findings as follows.
Foreseeing Danger and Constructing Defensive Structures:
As shown in Figure 7.a, AlphaStar (blue Zerg) had previously observed the Master Player’s (red Protoss) harassing air unit, Oracle, but failed to construct the anti-air Spore Crawler in advance for defense. This resulted in significant damage when the Master Player used Oracle for harassment, as AlphaStar had insufficient anti-air firepower and could only construct Spore Crawlers reactively. Conversely, the LLM agent, utilizing the Chain of Summarization method, can anticipate potential crises. As depicted, the LLM agent (green Protoss,Figure 7.b) constructed the defensive structure Shield Battery before the attack of the built-in AI (red Zerg), completing it just in time to assist the defense forces in successfully repelling the assault.
Adapting Unit Composition Based on Battlefield Situations:
AlphaStar fails to adjust its unit composition in response to the enemy’s forces, not utilizing the most effective troop configurations to defeat the opponent. Specifically, in Figure 13.a, AlphaStar (blue Zerg) uses numerous Queens to counter Terran’s Liberators, whereas it should have transitioned to producing Corruptors against Liberators. In Figure 13.b, AlphaStar (red Zerg) builds three Spore Crawlers to defend against the Master Player’s (blue Protoss) Shield Battery and Immortal rush. However, Spore Crawlers, being anti-air structures, offer minimal defense against Immortal attacks. In Figure 13.c, the Master Player’s (blue Protoss) army, composed of Immortals and Disruptors, perfectly counters the Zerg’s Roach forces. Yet, AlphaStar (red Zerg), despite being in a strong economic and technological position, continues to produce Roach units without transitioning. In Figure 13.d, the Master Player’s (red Terran) army employs a Bio-Tank (Marine, Medivacs, Marauders, Tanks) composition with almost no air units. AlphaStar (blue Zerg) opts for Corruptors, Ravagers, and Brood Lords, where the abundance of Corruptors is largely ineffective against the Master Player’s ground forces, also compromising mobility.
LLM agent adapts its unit composition based on the enemy’s forces, employing efficient troop configurations to defeat the opponent. In Figure 13.e, the LLM agent (green Protoss) uses Voidrays to counter the built-in AI’s (red Zerg) Roach and Hydra forces, continually producing Voidrays. Since the Roach and Hydra lack anti-air capabilities, Voidrays effectively handle them. In Figure 13.f, the LLM agent (red Protoss) chooses a Stalker and Immortal troop composition against the built-in AI’s (blue Zerg) Roach and Hydra forces. It also upgrades Psionic Storm technology, providing AoE damage effective against Roach and Hydra forces. In Figure 13.g, the built-in AI’s (blue Zerg) forces primarily consist of Ultralisks, while the LLM agent (red Protoss) is producing Immortals, countering Ultralisks through troop transition. In Figure 13.h, the built-in AI’s forces are composed of Banelings and Ultralisks. The LLM agent’s (red Protoss) troops are already capable of handling the opponent’s forces and are transitioning to air units, ensuring subsequent troop transitions and more diverse strategies.
LLM agent utilizes multi-frame summarization to analyze L1 results(Figure 14,16,18,20), achieving outputs(C,C,C,C) that are both interpretable and rational. Compared to AlphaStar, the decision-making(Figure 15,17,19,21) of the LLM agent is aided by human historical experience and logical reasoning. Consequently, in terms of interpretability and rationality, it surpasses AlphaStar.
6 Discussion
In this study, we successfully introduced TextStarCraft II, a textual representation of the starcraft II environment. This representation enables the LLM agent to interact with the starcraft II game solely through natural language. Our novel multi-level summarization approach allows the language model to understand the game’s progression and discern strategies employed by both allies and adversaries. Coupled with our tailored prompt design, the language model can leverage historical human gameplay experience and recognize in-game scenarios in starcraft II, subsequently deciding upon the most suitable strategies to adopt.
Compared to previous works on the starcraft II full game, our methodology stands out by achieving competitive results without the necessity for additional data and training. Specifically, by solely utilizing the gpt3.5 API, our prompt design, and the multi-level summarization technique, we demonstrated zero-shot capability in defeating the medium-hard built-in AI. Furthermore, our LLM agent exemplifies higher-order decision-making capabilities absent in prior works. It excels in Adaptive Strategic Adjustment(Fig6.a), aligning its strategies in real-time with the opponents’ troop transitions. The agent also demonstrates prowess in Technological Evolution and Innovation(Fig5.c), advancing in various tech domains proactively. It embodies Crisis Prevention and Management(Fig5.d) by constructing defensive structures preemptively, showcasing acute awareness of potential threats. Moreover, the agent’s Macro-economic and Defensive Strategizing(Fig5.c), coupled with its acumen in Intelligence Gathering and Analysis(Fig5.a and Fig5.b) illustrate its comprehensive strategic insight and versatile adaptability in StarCraftII’s intricate environments.
However, despite our achievements against several built-in AIs, we have identified several areas for improvement. Rule-Based Macro Actions: The current actions are heavily reliant on human prior knowledge as they are rule-based macro actions. We aim to mitigate this by incorporating hierarchical reinforcement learning in future implementations.Lack of Visual Information: The absence of visual input means the LLM agent lacks comprehensive battlefield awareness. Advancements in Visual Language Models (VLMs), such as miniGPT [33], are believed to have the potential to address this limitation.No Feedback Loop: The existing environment does not provide a feedback loop, rendering the multi-level LLM agent incapable of learning. The incorporation of a feedback mechanism is in our plans for future research.Memory Limitations: Due to API call speed constraints, our model only achieves two-tiered summarization, leading to a lack of both short-term and long-term memory in our LLM agent. This is another area we intend to explore in our subsequent work.
We posit that, moving forward, TextStarCraft II will serve as an excellent benchmark for assessing the capabilities of both LLMs and LLM agents. Pitting LLM agents against each other on TextStarCraft II offers a more intuitive and intricate evaluation approach than traditional dataset assessments. Moreover, we envision a fusion of LLM and reinforcement learning in future starcraft II full game AIs. Such a combination can potentially lead to enhanced strategic gameplay and a heightened degree of interpretability, setting a new standard beyond what has been achieved with AlphaStar.
Acknowledgments
We extend our sincere gratitude to the members of the starcraft II AI channel on the Discord application for their invaluable assistance and feedback. Special acknowledgment goes to the author of python-sc2, BuRny, whose generous support and guidance were instrumental in the completion of this work.
Appendix
Appendix A Prompt and data Example
@@
There are two types of system prompts in the 5.2 section.
@@
Appendix B Appendix: Questions and LLM Responses
@@
B.1 Basic Knowledge of StarCraft II
@@ 7. Free-to-Play Transition:
B.2 Racial Mechanics in StarCraft II
@@ 3. Terran:
B.3 Typical Strategies in StarCraft II for Different Matchups
@@
B.4 Ability to Outline Build Orders for Different Races
@@
Certainly, "StarCraft II" revolves heavily around build orders, which are pre-determined sets of actions aimed at maximizing efficiency, economy, and production. These build orders set the stage for your mid-game and late-game strategies. Below are standard build orders for each race, which are foundational and can be adapted based on opponents’ actions. Remember, the numbers represent the total worker count you should have when you begin each step. Terran (Standard Opening):
@@
Sure, here is a standard build order for each race in StarCraft II: Terran Bio Build Order:
@@
B.5 Classic Strategies and Counterplays in StarCraft II
@@
Appendix C Alphastar vs LLM Agent in interpretability
Example1: scouting probe
Figure 14: L1 results of 00:46 period
LLM output:
  • 3. Our Situation:
Suggestions:
  • Decisions:
    • 0: <BUILD NEXUS>
    • 1: <BUILD PYLON>
    • 2: <BUILD GATEWAY>
    • 3: <TRAIN PROBE>
    • 4: <SCOUTING PROBE>
Figure 15: Decision of LLM Agent: Scouting probe
Example2: chronoboost cyberneticscore
Figure 16: L1 results of 03:43 period
LLM output:
  • 3. Our Situation:
Suggestions:
  • Decisions:
    • 0: <RESEARCH WARPGATERESEARCH>
    • 1: <BUILD ZEALOT>
    • 2: <BUILD PROBE>
    • 3: <BUILD PYLON>
    • 4: <CHRONOBOOST CYBERNETICSCORE>
Figure 17: Decision of LLM Agent: Chronoboost Cybernetics Core
Example3: chronoboost nexus
Figure 18: L1 results of 04:25 period
LLM output:
  • 3. Our Situation:
Suggestions:
  • Decisions:
    • 0: <TRAIN PROBE>
    • 1: <BUILD GATEWAY>
    • 2: <EXPAND TO NEW RESOURCE LOCATION>
    • 3: <BUILD ROBOTICSFACILITY>
    • 4: <CHRONOBOOST NEXUS>
Figure 19: Decision of LLM Agent: Chronoboost Nexus
Example4: Adapt strategy in response to opponent’s tactics
Figure 20: L1 results of 13:28 period
LLM output:
  • 3. Our Situation:
    Suggestions:
    Decisions:
    • 0: <TRAIN PHOENIX>
    • 1: <TRAIN VOIDRAY>
    • 2: <BUILD STARGATE>
    • 3: <TRAIN STALKER>
    • 4: <TRAIN COLOSSUS>

### v2

@@ Appendix C Experiment Figure
Appendix D Prompt and data Example
@@
There are two types of system prompts in the 6.1 section.
@@
Appendix E Alphastar and llm agent
Appendix F Policy interpretability examples
In Section 6.3, we introduced the concept of policy interpretability within our LLM agent. We now present four illustrative examples to demonstrate how the LLM agent can generate reasonable strategies and decisions in TextStarCraft II. Each example comprises three components:
  • L1 Results Figure: This displays the input of the multi-frame summarization method.
  • LLM Output: The output of the multi-frame summarization method, which includes the LLM’s analysis, suggestions, and decisions.
  • Decision Figure: Illustrates the actual implementation within the StarCraft II game engine.
F.1 Example 1
Figure 14: L1 results of 00:46 period
LLM output:
  • 3. Our Situation:
Suggestions:
  • Decisions:
    • 0: <BUILD NEXUS>
    • 1: <BUILD PYLON>
    • 2: <BUILD GATEWAY>
    • 3: <TRAIN PROBE>
    • 4: <SCOUTING PROBE>
Figure 15: Decision of LLM Agent: Scouting probe
F.2 Example 2
Figure 16: L1 results of 03:43 period
LLM output:
  • 3. Our Situation:
Suggestions:
  • Decisions:
    • 0: <RESEARCH WARPGATERESEARCH>
    • 1: <BUILD ZEALOT>
    • 2: <BUILD PROBE>
    • 3: <BUILD PYLON>
    • 4: <CHRONOBOOST CYBERNETICSCORE>
Figure 17: Decision of LLM Agent: Chronoboost Cybernetics Core
F.3 Example 3
Figure 18: L1 results of 04:25 period
LLM output:
  • 3. Our Situation:
Suggestions:
  • Decisions:
    • 0: <TRAIN PROBE>
    • 1: <BUILD GATEWAY>
    • 2: <EXPAND TO NEW RESOURCE LOCATION>
    • 3: <BUILD ROBOTICSFACILITY>
    • 4: <CHRONOBOOST NEXUS>
Figure 19: Decision of LLM Agent: Chronoboost Nexus
F.4 Example 4
Figure 20: L1 results of 13:28 period
LLM output:
  • 3. Our Situation:
    Suggestions:
    Decisions:
    • 0: <TRAIN PHOENIX>
    • 1: <TRAIN VOIDRAY>
    • 2: <BUILD STARGATE>
    • 3: <TRAIN STALKER>
    • 4: <TRAIN COLOSSUS>
Appendix G Questions and LLM Responses
@@
G.1 Basic Knowledge of StarCraft II
@@ 7. Free-to-Play Transition:
G.2 Racial Mechanics in StarCraft II
@@ 3. Terran:
G.3 Typical Strategies in StarCraft II for Different Matchups
@@
G.4 Ability to Outline Build Orders for Different Races
@@
Certainly, "StarCraft II" revolves heavily around build orders, which are pre-determined sets of actions aimed at maximizing efficiency, economy, and production. These build orders set the stage for your mid-game and late-game strategies. Below are standard build orders for each race, which are foundational and can be adapted based on opponents’ actions. Remember, the numbers represent the total worker count you should have when you begin each step.
Terran (Standard Opening):
@@
Sure, here is a standard build order for each race in StarCraft II: Terran
Bio Build Order:
@@
G.5 Classic Strategies and Counterplays in StarCraft II
@@
Appendix H NeurIPS Paper Checklist
H.1 Claims
Answer: [Yes]
Justification: The abstract and introduction clearly outline the main claims and contributions of developing a new method for integrating LLMs into RTS gaming, particularly StarCraft II, and improving model interpretability. These sections accurately reflect the scope of the research and its applicability to other complex decision-making environments
H.2 Limitations
Answer: [Yes]
Justification: TextStarCraft II’s limitations include its reliance on non-visual data inputs, the use of scripted micro policies, and constrained performance due to resource limitations. For more detailed information, please refer to the Discussion section 7.
H.3 Theory Assumptions and Proofs
Answer: [NA]
Justification: The paper primarily focuses on experimental results and method development rather than theoretical proofs.
H.4 Experimental Result Reproducibility
Answer: [Yes]
Justification: The paper details all necessary configurations, methodologies, and the TextStarCraft II environment used, ensuring reproducibility. Additionally, we have open-sourced the code, data, and a demo, providing full access for researchers to replicate and verify our results comprehensively.
H.5 Open access to data and code
Answer: [Yes]
Justification:We have provided open access to our complete source code and dataset, ensuring full reproducibility of our experiments. All resources, including training and testing scripts, are available in a publicly accessible repository linked in the supplemental materials of our paper. We also provide the anonymous link:https://anonymous.4open.science/r/Large-Language-Models-play-StarCraftII-8C45/readme.md.
H.6 Experimental Setting/Details
Answer: [Yes]
Justification: Our experimental setup is thoroughly detailed in Section B.1, including comprehensive descriptions of agent and opponent selection, map choice, parameter settings, and more, ensuring a robust and reproducible experimental environment.
H.7 Experiment Statistical Significance
Answer: [Yes]
Justification: Our paper presents error bars calculated as the standard deviation across multiple experiment runs, capturing variability due to initialization. These error bars are clearly displayed and annotated in the experimental result tables and figures.
H.8 Experiments Compute Resources
Answer: [Yes]
Justification: we have introduced our experiments compute resources, detailed in B.3.
H.9 Code Of Ethics
Answer: [Yes]
Justification: Our research complies with the NeurIPS Code of Ethics, ensuring ethical standards and transparency in all experimental procedures, with measures in place to mitigate any potential risks, detailed in Code Of Ethics.
H.10 Broader Impacts
Answer: [Yes]
Justification:The paper discusses both positive and negative societal impacts of using LLMs in strategic environments, detailed in Broader Impact.
H.11 Safeguards
Answer: [NA]
Justification: Our research does not introduce data or models with a high risk for misuse. The project focuses on developing strategic AI capabilities within the controlled environment of StarCraft II using TextStarCraft II, which poses no inherent risk of dual-use or misuse. Thus, no specific safeguards beyond standard ethical research practices were required.
H.12 Licenses for existing assets
Answer: [Yes]
Justification: All used assets are properly credited with detailed references and URLs. We explicitly state each asset’s license type, predominantly CC-BY 4.0, ensuring all terms of use are respected and documented as per compliance standards.
H.13 New Assets
Answer: [Yes]
Justification: We introduced TextStarCraft II, a new, thoroughly documented asset. Documentation including setup, usage, and licensing details is available in a publicly accessible, anonymized repository, ensuring transparency and integrity in the review process. https://anonymous.4open.science/r/Large-Language-Models-play-StarCraftII-8C45/readme.md
H.14 Crowdsourcing and Research with Human Subjects
Answer: [Yes]
Justification: We involved 30 StarCraft II masters and grandmasters, including professional and semi-professional players, in our study. For more details on the participant engagement and data handling, please see the section referenced in B.4.
H.15 Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
Answer: [NO]
Justification:Although our study involved human participants, their engagement was restricted to game-related QA and gameplay testing, which are low-risk activities typical of gaming studies. This type of interaction generally does not require IRB approval as it does not involve exposure to physical or psychological risks beyond those encountered in normal gameplay. Participants were fully briefed on the nature of the tests and their voluntary participation was secured, aligning with ethical standards for such research.